OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 2x + y = 6

OpenStudy (anonymous):

6 is c in this equation. when you subtract it from both sides you set the equation equal to 0. that should help you find the value of c

OpenStudy (anonymous):

2x + y + 6 = 0 ?

jimthompson5910 (jim_thompson5910):

close, more like 2x + y + (-6) = 0 or 2x + y - 6 = 0

jimthompson5910 (jim_thompson5910):

2x + y + (-6) = 0 matches up with Ax + By + C = 0 A = 2 B = 1 C = -6